# Beam Enumeration: Probabilistic Explainability For Sample Efficient Self-conditioned Molecular Design

### v1

@@ D.4 Wall Times
The wall times for all drug discovery case studies with every algorithm is presented in Table 12. The reported values are averaged over 3 replicates. In general, adding Beam Enumeration to the base Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a) algorithms increased wall times but only slightly and it is negligible when considering expensive oracles. An interesting observation is that "sample" token sampling increases wall time variance. This is because less probable substructures lead to more "filter rounds’, i.e., epochs where all the sampled molecules are discarded as none of them contain the Beam Enumeration extracted substructures. In addition, REINVENT generally has longer wall times even though the oracle budget is the same. The reason for this is because REINVENT optimizes the structure components of the MPO objective: QEDBickerton et al. (2012) and MW constraint to a lesser extent. Consequently, REINVENT generates larger molecules, on average, which take longer to dock with VinaTrott and Olson (2010). This observation is in agreement with the original Augmented Memory work which compared to REINVENT.
Appendix E Augmented Memory and REINVENT Model Hyperparameters
Table 13: LSTM model hyperparameters for Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a)
@@ Appendix E Augmented Memory and REINVENT Model Hyperparameters
The same pre-trained prior on ChEMBLGaulton et al. (2012b) was used for Augmented MemoryGuo and Schwaller (2023) and REINVENTOlivecrona et al. (2017); Blaschke et al. (2020a). All shared hyperparameters (sampling batch size and learning rate) are the same. Default additional hyperparameters for Augmented Memory were used based on the original workGuo and Schwaller (2023): two augmentation rounds and using Selective Memory Purge to prevent mode collapse. Experience replayLin ; Blaschke et al. (2020a) was kept default in REINVENT (randomly sample 10 molecules out of 100 from the replay buffer at each epoch).

### v2

@@ E.4 Wall Times
The wall times for all drug discovery case studies with every algorithm is presented in Table 12. The reported values are averaged over 3 replicates. In general, adding Beam Enumeration to the base Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020) algorithms increased wall times but only slightly and it is negligible when considering expensive oracles. An interesting observation is that ”sample” token sampling increases wall time variance. This is because less probable substructures lead to more ”filter rounds’, i.e., epochs where all the sampled molecules are discarded as none of them contain the Beam Enumeration extracted substructures. In addition, REINVENT generally has longer wall times even though the oracle budget is the same. The reason for this is because REINVENT optimizes the structure components of the MPO objective: QED (Bickerton et al. (2012) and MW constraint to a lesser extent. Consequently, REINVENT generates larger molecules, on average, which take longer to dock with Vina (Trott & Olson (2010). This observation is in agreement with the original Augmented Memory work which compared to REINVENT.
Appendix F Augmented Memory and REINVENT Model Hyperparameters
Table 13: LSTM model hyperparameters for Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020)
@@ Appendix F Augmented Memory and REINVENT Model Hyperparameters
The same pre-trained prior on ChEMBL (Gaulton et al. (2012b) was used for Augmented Memory (Guo & Schwaller (2023) and REINVENT (Olivecrona et al. (2017); Blaschke et al. (2020). All shared hyperparameters (sampling batch size and learning rate) are the same. Default additional hyperparameters for Augmented Memory were used based on the original work (Guo & Schwaller (2023): two augmentation rounds and using Selective Memory Purge to prevent mode collapse. Experience replay (Lin (1992); Blaschke et al. (2020) was kept default in REINVENT (randomly sample 10 molecules out of 100 from the replay buffer at each epoch).
Appendix G Beam Enumeration: Beyond Language-based Formulations
This section discusses potential extensions of Beam Enumeration to other classes of molecular generative models beyond language-based formulations.
G.1 Direct Application
Beam Enumeration enumerates from local probability distributions and is thus well-suited for autoregressive models. Correspondingly, autoregressive molecular generative models are not limited to language-based formulations. As a concrete example, GraphINVENT (Mercado et al. (2021b; a) generates molecular graphs and has been coupled with RL (Atance et al. (2022). During the optimization process, instead of token probabilities being updated, graph action probabilities are tuned, which denote discrete actions to construct the molecular graph, such as adding a node or edge. Beam Enumeration can be directly applied by exhaustively enumerating the most probable set of graph actions and extracting substructures (conditional on the molecular graph being valid).
G.2 Applying Observed Findings
Many generative models construct a substructure ”vocabulary” such that the optimization process is defined by how to combine these building blocks (Jin et al. (2018); Guo et al. (2022). This is similar to token sampling, except the ”tokens” are now ”words” (groups of atoms). The empirical results from Beam Enumeration show that substructure biasing is beneficial, even early in a generative experiment when the model has not necessarily learned to generate favorable molecules. Given a promising seed molecule, existing models such as the JT-VAE (Jin et al. (2018) and the Data Efficient Graph Grammar (DEG) (Guo et al. (2022) models can be used to propose local neighbourhood expansions. Scoring these molecules enable the possibility of iteratively performing substructure (of increasing size) biasing, effectively leveraging both the advantages of the learned ”vocabulary” of building blocks, and the iterative construction of increasingly favorable substructures. Results in this work suggest that this can be performant as enforcing larger substructure size always improved sample efficiency (up to a size of 15 in this work).
Appendix H Beam Enumeration vs. Beam Search
Beam Enumeration is distinct from Beam Search (Graves (2012); Boulanger-Lewandowski et al. (2013). In this section, we contrast the behavior point-by-point.
Objective: Beam Search approximates finding the highest probability sequences by maintaining a ”width” of the top k candidates at each step. The number of sequences decoded via Beam Search is exactly k. By contrast, Beam Enumeration exhaustively enumerates the top k tokens at each step to extract substructures, such that the total number of sub-sequences is $k^{Beam\>Steps}$.
Search Space: Beam Search keeps the top k candidates globally, pruning less probable sequences. Beam Enumeration keeps the top k local candidates per sub-sequence.
Termination: Beam Search terminates when a sequence ends or a maximum length is reached. Beam enumeration does a pre-defined number of steps, controlled by Beam Steps.
Use of results: Beam Search returns the highest probability candidates for direct use (for instance, as full decoded molecules in Moret et al. (2021). Beam Enumeration uses the enumerated sub-sequences to extract molecular substructures for self-conditioned generation.
In summary, Beam Enumeration performs local, exhaustive enumeration of sub-sequences to extract meaningful substructures, while Beam Search approximates finding a set of globally optimal (by probability) sequences. The exhaustive sub-sequence enumeration enables probabilistic extraction of knowledge from the model.